The Magazine Mass Marketing Company has received 15 entries in its latest sweepstakes. They know that the probability of receiving a magazine subscription order with an entry form is 0.5. What is the probability that exactly 2 of the entry forms will include an order? Round your answer to four decimal places.
Answer #1

Answer: Criven no 15 X = 2 P:0.5 I-D=1-0.5 =0.5 using binomial potability formule PLX=x) = Q paci-P]n-4 P(x+2) = (3) 60-53 (0